Using the appropriate interest table, answer the following questions. (Each case is independent of the others) Click here to view factor tables what is the future value of 24 periodic payments of $4,130 each made at the beginning of each period an to O decimal places, eg. 458,581.) Round actor values to 5 dec compounded at 8% a places, eg and nai answer The future value eTextbook and Media Click here to view factor tables What is the present value of $3,290 to be received at the beginning of each of 27 periods, discounted at 5% compound interest? Round actor values to 5 decimal places, eg. 1.25124 and final answer to O decimal places, eg. 458,581.) $ The present value
Click here to view factor tables What is the future value of 16 deposits。 3,540 each made at the beginning of each period and compounded at 10%? (Future value as of the end of the 16th period.) (Round factor values to 5 decimal places, eg. 1.25124 and final answer to O decimal places, eg. 458,581.) The future value eTextbook and Media Click here to view factor tables What is the present value of 8 receipts of $3.160 each received at the beginning of each period, discounted at 9% compounded interest? Round factor values to 5 decimal places, eg 1.25 124 and inal answer to O decimal places, eg. 458,581.) The present value
Answer #1

a)

=FV(8%,24,-4130,,1)

=297797.53

b)

=PV(5%,27,-3290,,1)

=50584.36

c)

=FV(10%,16,-3540,,1)

=139988.25

d)

=PV(9%,8,-3160)

=17490.03